OpenStudy (anonymous):

find the line tangent to y= 10 sqrt x - x+3 at the points (100,3)

hartnn (hartnn):

could you first get the derivative of y ? because we need the slope to find the equation, and slope = dy/dx at that time.

OpenStudy (anonymous):

wouldn't you use the y1-y2 = M( x1 - x2) equation?

hartnn (hartnn):

ofcourse, but we need to find M = slope, right ? \(\Large M= \dfrac{dy}{dx}|_{x=100}\)

hartnn (hartnn):

so first find dy/dx :) can u ?

OpenStudy (anonymous):

i forget how to find the derivative of this one but i can try...

hartnn (hartnn):

god :) if you get stuck, let me know, i'll help :)

hartnn (hartnn):

**good

OpenStudy (anonymous):

so i tried solving for y i plugged in 100 for x and got y=3 so would that be my y1?

hartnn (hartnn):

do you know what derivative is ?

hartnn (hartnn):

\(\Large \dfrac{d}{dx}x^n=n \times x^{n-1}\) seen anything like this before ?

OpenStudy (anonymous):

no not like that i'm sorry

hartnn (hartnn):

(then how are you attempting a derivative Question, when it is not taught to you ?) sqrt x = x^(1/2) so put n =1/2 in that formula to get derivative of sqrt x first.

OpenStudy (anonymous):

\[y = 10\sqrt{x} - x + 3 when (100,3) \]

OpenStudy (anonymous):

so you want me to find the derivatives of the inside

OpenStudy (anonymous):

soo would it be 5x^(-1/2) -1 ?

hartnn (hartnn):

correct! :) good. now just put x= 100 there! and you will get the value of M

hartnn (hartnn):

`slope of tangent to a curve at a point= derivative of the curve at that point`

OpenStudy (anonymous):

so M = 22.360679775?

hartnn (hartnn):

jow :O which calculator did u use ? \(\large M = 5 \times 100^{-1/2}-1 = 5 \times \dfrac{1}{\sqrt{100}}-1=5/10 -1=-1/2 \) got this ?

OpenStudy (anonymous):

oh jeez yes you're rihgt

hartnn (hartnn):

so we have M =-1/2 x1 =100, y1 =3 just plug these in y1-y2 = M( x1 - x2) and you'll get the equation of line :)

OpenStudy (anonymous):

3 - y^2 = 1/2(100 -x^2) don't i need another variable?

hartnn (hartnn):

how did u get y^1 and x^2 ?

hartnn (hartnn):